OpenStudy (elise_a18):

Simplify the square root of 189. - the square root of 21. + the square root of 84

OpenStudy (anonymous):

i think you get everything in terms of the square root of 21

OpenStudy (anonymous):

189 = 13.74772708

OpenStudy (elise_a18):

How though? @satellite73

OpenStudy (k_lynn):

You want to figure this out, @Elise_a18 \[\sqrt{189}-\sqrt{21}+\sqrt{84}\]?

OpenStudy (elise_a18):

correct :)

OpenStudy (elise_a18):

@k_lynn

OpenStudy (elise_a18):

nope

OpenStudy (elise_a18):

i just need to simplify the radicals, not solve

OpenStudy (elise_a18):

@k_lynn @Kainui @sammixboo @sleepyhead314

OpenStudy (sleepyhead314):

189 is 9 times 21 84 is 4 times 21

sammixboo (sammixboo):

Ok, so we have \(\color{blueviolet}{\sqrt{189}-\sqrt{21}+\sqrt{84}}\) Now, let me do the next step for ya \(\color{blueviolet}{\sqrt{189}-\sqrt{21}+\sqrt{84}}\) \(\color{blueviolet}{\sqrt{21\times9}-\sqrt{21}+\sqrt{21\times4}}\) \(\color{blueviolet}{\sqrt{21}\sqrt{9}-\sqrt{21}+\sqrt{21}\sqrt{4}}\) Now, what is the square root of 9 and 4?

sammixboo (sammixboo):

I am awful an explaining how to solve radicals, so I am trying my best Dx

OpenStudy (sleepyhead314):

:O PURTY LATEX

sammixboo (sammixboo):

I am making sure what I am doing is correct also, so it might take me a little extra time

OpenStudy (elise_a18):

You're doing fine Sammiiiii :3 ................3 and 2

OpenStudy (elise_a18):

Otay! :D

sammixboo (sammixboo):

Right! So \(\color{blueviolet}{\sqrt{21}\sqrt{9}-\sqrt{21}+\sqrt{21}\sqrt{4}}\) \(\color{blueviolet}{3\sqrt{21}-\sqrt{21}+2\sqrt{21}}\) **remember \(\color{blueviolet}{\sqrt{21}}\) is the same as \(\color{blueviolet}{1\sqrt{21}}\), so we are going to use that** \(\color{blueviolet}{3\sqrt{21}-1\sqrt{21}+2\sqrt{21}}\)

sammixboo (sammixboo):

Let me refresh real quick. Highlighting isn't working D:

sammixboo (sammixboo):

OK back

sammixboo (sammixboo):

Now after we have \(\color{blueviolet}{3\sqrt{21}-1\sqrt{21}+2\sqrt{21}}\) Let's look at the numbers that are not a radical. \(\color{blueviolet}{3\color{black}{\sqrt{21}}-1\color{black}{\sqrt{21}}+2\color{black}{\sqrt{21}}}\) \(\color{blueviolet}{3-1+2}\) What's 3 - 1 + 2?

OpenStudy (elise_a18):

So it's be \[\sqrt[4]{21}\] ? :D

sammixboo (sammixboo):

Almost! Just got the 4 in the wrong place. So 3 - 1 + 2 = 4. After that, you just plop the four right infront of the radical, because you are multiplying it. It isn't going ot be a pat of the radical though, so you would have \(\color{blueviolet}{4\sqrt{21}}\)

sammixboo (sammixboo):

I haven't done radicals in a long time hehaha e-e I forgot what the different number placements are called in a radical

OpenStudy (nincompoop):

\(\sqrt[4]{21} \not= 4\sqrt{21} \)

sammixboo (sammixboo):

Yeah, understand Elise :P I guess if you need anything else let me know